# FireANTs: Adaptive Riemannian Optimization for Multi-Scale Diffeomorphic Registration

### v1

@@ 4 Methods
Given $d$-dimensional images $I:\Omega\to\mathbb{R}^{d}$ and $I^{\prime}:\Omega\to\mathbb{R}^{d}$ where the domain $\Omega$ is a compact subset of $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$, image registration is formulated as an optimization problem to find a transformation $\varphi$ that warps $I^{\prime}$ to $I$. The transformation can belong to a group, say $G$, whose elements $g\in G$ act on the image by transforming the domain as $(I\circ g)(x)=I(g(x))$ for all $x\in\Omega$. The registration problem solves for
@@
where $C$ is a cost function, e.g., that matches the pixel intensities of the warped image with those of the fixed image, or local normalized cross-correlation or mutual information across image patches. There are many types of regularizers $R$ used in practice, e.g., total variation, elastic regularization 33, enforcing the transformation to be invertible 34, or volume-preserving 76 using constraints on the determinant of the Jacobian of $\varphi$, etc. If, in addition to the pixel intensities, one also has access to label maps or different anatomical regions marked with correspondences across the two images, the cost $C$ can be modified to ensure that $\varphi$ transforms these label maps or landmarks appropriately.
@@
We perform registration over the group of diffeomorphisms $G=\text{Diff}(\Omega)$; a diffeomorphism is a smooth and invertible map with a corresponding differentiable inverse map 77, 78, 79. It is useful to note that unlike rigid or affine transforms that have a fixed number of parameters, the number of parameters in a diffeomorphism scales with the size of the domain. When groups of transformations on continuous domains are endowed with a differentiable structure, they are called Lie groups. Lie groups equipped with a Riemannian metric are Riemannian manifolds.
To summarize, there are three key parts of registration methods: the objective, the group $G$, and the optimization algorithm. Our work focuses on developing new optimization algorithms.
Euclidean gradient descent using the Lie algebra in shooting methods
Each Lie group has a corresponding Lie algebra $\mathfrak{g}$ which is the tangent space at identity. This creates a one-to-one correspondence between elements of the group $g\in G$ and elements of its Lie algebra $v\in\mathfrak{g}$ given by the exponential map
$\exp:\mathfrak{g}\to G;$
@@
effectively to reach $g=\exp(v)\text{Id}$ from identity $\text{Id}\in G$, the exponential map says that we have to move along $v$ for unit time. Exponential maps for many groups can be computed analytically, e.g., Rodrigues transformation for rotations, Jordan-Chevalley decomposition 80, or the Cayley Hamilton theorem 81 for matrices. For diffeomorphisms, the Lie algebra is the space of all smooth velocity fields $v:\Omega\to\mathbb{R}^{d}$. There exist iterative methods to approximate the exponential map called the scaling-and-squaring approach 53, 46 which uses the identity
@@
By virtue of the exponential map, we can solve the registration problem of finding $\varphi\in G$ by directly optimizing over the Lie algebra $v$. This is because the Lie algebra is a vector space and we can perform, for example, standard Euclidean gradient descent for registration 82, 83, 84. Such methods are called stationary velocity field or shooting methods. At each iteration, one uses the exponential map to get the transformation $\varphi$ from the velocity field $v$, computes the gradient of the registration objective with respect to $\varphi$, pulls back this gradient into the tangent space where $v$ lies
$\nabla_{v}L=\partialderivative{\varphi}{v}\nabla_{\varphi}L$
and finally makes an update to $v$. Traditional methods like DARTEL 53 implement this approach. This is also very commonly used by deep learning methods for registration 46, 8, 42 due to its simplicity. Geodesic shooting methods are more sophisticated implementations of this approach where the diffeomorphism $\varphi$ is the solution of a time-dependent velocity which follows the geodesic equation; the geodesic is completely determined by the initial velocity $v_{0}\in\mathfrak{g}$.
Riemannian gradient descent
Solving the registration problem directly on the space of diffeomorphisms avoids repeated computations to and fro via the exponential map. The downside however is that one now has to explicitly account for the curvature and tangent spaces of the manifold. The updates for Riemannian gradient descent 59 at the $t^{\text{th}}$ iteration are
@@
$\nabla_{\varphi}L$ $=\mathbf{g}^{-1}_{\varphi_{t}}\frac{\partial L}{\partial\varphi},$
@@
where one pulls back the Euclidean gradient $\frac{\partial L}{\partial\varphi}$ onto the manifold using the inverse metric tensor $\mathbf{g}$ (which makes the gradient invariant to the parameterization of the manifold of diffeomorphisms) before projecting it to the tangent space using $\text{Proj}_{\varphi_{t}}$. Since the tangent space is a local first-order approximation of the manifold’s surface, we can move along this descent direction by a step-size $\eta$ and compute the updated diffeomorphism $\varphi_{t+1}$, represented as the exponential map from $\varphi_{t}$ computed in the direction of $-\text{Proj}_{\varphi_{t}}(\nabla_{\varphi}L)$. In our work, we take advantage of a few key properties of diffeomorphisms. (a) If the step-size $\eta$ is small, the exponential map can be well approximated with a retraction map, which is quick to compute  44, 45. (b) The metric tensor is the Jacobian of the diffeomorphism $\varphi_{t}$ 79, 40 which can be approximated with finite differences. (c) The tangent space is the set of all $C^{\infty}$ velocity fields, which is the same as the ambient space of the diffeomorphisms and therefore we can omit the projection step. We also implement a stochastic variant of Riemannian gradient descent whereby we only update the diffeomorphism on a subset of the image domain; convergence properties of this algorithm can be studied 57.
For high-dimensional groups like diffeomorphisms, optimizing the transformation directly on the manifold is preferable to optimizing on the Lie algebra. We noticed this empirically in a number of instances. In Fig. 3, the greedy SyN method (which performs Riemannian optimization) outperforms the Lie algebra variant (DARTEL) significantly, runs much faster on average (computation of the exponential map and its derivative adds additional time and memory overhead), and results in substantially fewer singularities in the velocity field. Similar observations may be made for EMPIRE10 dataset in the ANTs baseline. In  Fig. S.6 we observed the across a large variety of hyper-parameters (obtained via grid search), Riemannian gradient descent leads to better target overlap compared to the Lie algebra variant on the LPBA40 dataset.
Adaptive Riemannian optimization for diffeomorphisms
Adaptive optimization algorithms such as RMSProp 54, Adagrad 56 and Adam 55 have become popular because they can handle poorly conditioned optimization problems in deep learning. Variants for optimization on low-dimensional Riemannian manifold exist 57, 85, 86, 58. Diffeomorphisms are a high-dimensional group (e.g., the size of velocity field scales with that of the domain). Also, often the number of parameters (e.g., size of the image) in these methods is fixed which makes it difficult to run them on diverse datasets and modalities. We develop a multi-scale 60, 87, 45, 53 approach to optimization on Riemannian manifolds that can adapt the updates to the curvature of the manifold and that work for pairs of images of different sizes.
Adaptive optimization methods, in Euclidean space, typically maintain a moving average of past gradients (momentum) and an approximation of the Hessian (which allows approximate second-order updates). The Hessian is generally expensive to compute and store, and therefore only diagonal elements are sometimes computed; one may resort to further approximations (like Adam does) and maintain a running average of the element-wise squared gradients (we will call this the “curvature vector”). Both the momentum and the curvature vector can be thought of as vectors in the tangent space. For Euclidean manifolds, the tangent space is the same as the manifold and it is easy to compute the modified descent direction by transporting the momentum and the curvature vectors along a straight line; in Euclidean space such transport does not change the magnitude or direction of a vector. On curved manifolds, parallel transport generalizes the notion of transporting a vector $v$ from the paths connecting a point $\varphi$ to another $\varphi^{\prime}$. And unlike Euclidean space, parallel transport depends upon the path between the two points. It is expensive to compute parallel transport for groups such as diffeomorphisms. This makes it difficult and expensive to implement adaptive optimization methods.
We can work around the above issue using a result of Younes et al. 79 on computing the differentials at any transformation $\varphi$ using the differential at identity Id. Let us first rewrite Eq. 2 in a slightly different notation. The Eulerian differential ¯$\bar{\partial}L(\varphi)$ is a linear map (also called a linear form) from vector fields on $\Omega$ to real numbers, and denotes the change in $L$ when $\varphi$ is changed along a velocity field $v$
(¯$\left(\bar{\partial}L(\varphi)\mid v\right)=\partial_{\epsilon}L(\varphi+\epsilon(v\circ\varphi))|_{\epsilon=0}.$
Much like standard gradient descent in Euclidean space, iterative updates to the diffeomorphism $\varphi$ using the Eulerian differential minimize the objective $L$. We have
(¯$\left(\bar{\partial}L(\varphi)\mid v\right)=\left(\frac{\delta L}{\delta\varphi}\mid v\circ\varphi\right),$
for any velocity field $v$. This is a direct correspondence between the Eulerian differential that performs Riemannian gradient descent in Eq. 2 on the left-hand side and the conventional derivative that can be calculated analytically on the right-hand side. Exploiting this correspondence for optimization requires computing $v\circ\varphi$ each time. But Younes et al.show in Section 10.2 of their book 79 that:
(¯$\left(\bar{\partial}L(\varphi;\ I,I^{\prime})\mid v\right)=\left(\bar{\partial}L(\text{Id};\ I,I^{\prime}\circ\varphi)\mid v\right).$ (3)
This allows us to represent the Riemannian gradient at arbitrary $\varphi$ (left) in terms of the gradient at $\varphi=\text{Id}$ calculated for the deformed image $I^{\prime}\circ\varphi$ (right). In simpler words, we can pretend as if the optimization algorithm always works at identity Id at every iteration if we match to a warped image $I^{\prime}\circ\varphi_{t}$. When Riemannian gradient descent is implemented like this, gradients, momentum and curvature vector lie in the tangent space at identity for all iterations, and calculating the gradient descent update is therefore identical to that of the Euclidean case. Parallel transport is not required. The Riemannian metric tensor $\mathbf{g}_{\varphi_{t}}$ is also the outer product of the Jacobian of the diffeomorphism at identity; this is identity. We therefore do not need to pullback the gradient in Eq. 2 on the manifold. This is a very useful technique that eliminates a number of computationally expensive steps. We should emphasize that it is mathematically rigorous and does not result from any approximations. We illustrate this procedure in  Fig. S.3(a).
Interpolation strategies for multi-scale registration

### v2

@@ 4 Methods
4.1 Preliminaries
Given $d$-dimensional images $I:\Omega\to\mathbb{R}^{d}$ and $I^{\prime}:\Omega\to\mathbb{R}^{d}$ where the domain $\Omega$ is a compact subset of $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$, image registration is formulated as an optimization problem to find a transformation $\varphi$ that warps $I^{\prime}$ to $I$. The transformation can belong to an algebraic group, say $G$, whose elements $g\in G$ act on the image by transforming the domain as $(I\circ g)(x)=I(g(x))$ for all $x\in\Omega$. The registration problem solves for
@@ 4.1 Preliminaries
where $C$ is a cost function, e.g., that matches the pixel intensities of the warped image with those of the fixed image, or local normalized cross-correlation or mutual information of image patches. There are many types of regularizers $R$ used in practice, e.g., total variation, elastic regularization 49, enforcing the transformation to be invertible 50, or volume-preserving 51 using constraints on the determinant of the Jacobian of $\varphi$, etc. If, in addition to the pixel intensities, one also has access to label maps or different anatomical regions marked with correspondences across the two images, the cost $C$ can be modified to ensure that $\varphi$ transforms these label maps or landmarks appropriately.
4.1.1 Properties of the considered Transformation Group
@@
A diffeomorphism is a smooth and invertible map with a corresponding differentiable inverse map 52, 53, 54. We denote the set of all diffeomorphisms on $\Omega$ as $\text{Diff}(\Omega;\mathbb{R}^{d})$. It is useful to note that unlike rigid or affine transforms that have a fixed number of parameters, diffeomorphisms require dense and variable parameterization, typically proportional to the size of the image. When groups of transformations on continuous domains are endowed with a differentiable structure, they are called Lie groups. Diffeomorphisms are also examples of Riemannian manifolds, and are amenable to Riemannian optimization (see  Section 4.3.2).
In this work, we only consider a subgroup of diffeomorphisms. Consider the set of continuously differentiable functions $u\in C^{1}_{0}(\Omega,\mathbb{R}^{d})$ such that $u,J(u)=0$ on $\partial\Omega$, where $J(u)$ is the Jacobian of $u$, such that $[J(u)(x)]_{ij}=\frac{\partial u(x)_{i}}{\partial x_{j}}$. These functions can be extended to have $u\equiv 0$ outside $\Omega$. Then, for a small enough $\epsilon>0$, $x+\epsilon u(x)$ is a diffeomorphism (Proof in Proposition 8.6 in  54). Although these diffeomorphisms are close to identity, diffeomorphisms with larger deviations from the identity can be constructed by composing these ‘small diffeomorphisms’. Therefore, we study the subgroup of diffeomorphisms of the form
$\phi_{n}=(id+\epsilon_{1}u_{1})\circ\ldots\circ(id+\epsilon_{n}u_{n})$ (6)
where $u_{i}$s are defined as before. We denote this subgroup as $G(\Omega,\mathbb{R}^{d})$. This subgroup retains the group structure with identity element $id$, the composition operation $\circ$ induced from $\text{Diff}(\Omega,\mathbb{R}^{d})$, and the inverse group element: $\phi_{n}^{(-1)}=(id+\epsilon_{n}u_{n})^{(-1)}\circ\ldots\circ(id+\epsilon_{1}u_{1})^{(-1)}$ (as each individual $id+\epsilon_{n}u_{n}$ is shown to have an inverse 54). The elements of this subgroup can be thought of as diffeomorphisms arising from time-varying continuously differentiable flows.
However, the rate of convergence of these algorithms are contingent on the severity of ill-conditioning of  Eq. 5. In subsequent sections, we first show the extent of ill-conditioning for diffeomorphic registration which subsequently warrants adaptive optimization over this subgroup of diffeomorphisms.
4.2 Deformable Image Registration is a severely ill-conditioned problem
The ill-conditioned nature of image registration represents a comparatively neglected domain of inquiry within the extant literature. Recent works in the literature  55, 56 only speculate the ill-conditioned nature of registration but do not quantify it. Computing the ill-conditioning requires us to analyze the Hessian of the registration cost function. This is infeasible in general due to the high dimensionality of the problem; the full Hessian of a MRI brain registration problem requires more than 15 petabytes of memory to store. However, we consider a typical scenario of T1-weighted 3D MRI image registration with the L2 loss 30, 15, 16: i.e. $C(I_{f},I_{m},\varphi)=\sum_{i}(I_{f}(\mathbf{x}_{i})-I_{m}(\varphi(\mathbf{x}_{i})))^{2}$. In this case, the gradient of $C$ w.r.t. $\varphi(\mathbf{x}_{i})$ is $(I_{m}(\varphi(\mathbf{x}_{i}))-I_{f}(\mathbf{x}_{i}))\nabla I_{m}(\varphi(\mathbf{x}_{i}))$, which does not depend on $\varphi(\mathbf{x}_{j}),j\neq i$. Therefore, the full Hessian is simply a block-diagonal matrix containing pixelwise Hessians $H_{i}=\nabla^{2}_{\varphi(\mathbf{x}_{i})}C$ with eigenvalues $\{\lambda_{i};i=\{1,2,3\}\}$. This makes the conditioning analysis tractable. We calculate the per-pixel condition number, defined as $\kappa_{i}=|\lambda_{i}^{\text{max}}|/|\lambda_{i}^{\text{min}}|$; and investigate the relationship between the fraction of foreground pixels and $\kappa_{i}$ across multiple spatial resolutions of the images. The study considers three downsampling factors: 1x (original resolution), 2x, and 4x, in accordance with existing multi-scale optimization techniques. Fig. 6 shows that across all resolutions, more than 60% of foreground pixels have a condition number greater than 10. To elucidate the impact of poor conditioning on optimization, we construct a simplified example of an ill-conditioned two-dimensional convex optimization problem, detailed in Section A.3. Even with $\kappa=10$, convergence slows down drastically for an ill-conditioned convex optimization problem. This indicates severe ill-conditioning of the registration problem, strongly motivating the need for first-order adaptive optimization.
Figure 6: Deformable image registration is ill-conditioned. To quantitatively examine ill-conditioning in registration, we compute the distribution of per-pixel condition number for a MRI registration task, at different image downsampling factors (denoted as 1x, 2x, and 4x). A high condition number signifies exacerbated ill conditioning and requires higher-order optimization. A horizontal dashed line denoting $\kappa=10$ is drawn as a reference for substantial ill conditioning. Across all scales, a substantial fraction of foreground voxels are ill-conditioned ($\kappa>10$), necessitating adaptive first-order optimization for faster convergence and accurate registration.
4.3 Adaptive Optimization for Diffeomorphisms
We provide a brief overview of the mathematical frameworks employed to optimize parameters that reside on Riemannian manifolds like diffeomorphisms, followed by a novel algorithm that exploits the group action to define a gradient descent algorithm that eliminates computationally expensive steps. This novel formulation of the ‘gradient descent’ algorithm can then be formulated to incorporate adaptive algorithms such as Adam 26 to optimize diffeomorphisms.
4.3.1 Euclidean gradient descent using the Lie algebra in shooting methods
@@
Each Lie group has a corresponding Lie algebra $\mathfrak{g}$ which is the tangent space at identity. This creates a locally one-to-one correspondence between elements of the group $g\in G$ and elements of its Lie algebra $v\in\mathfrak{g}$ given by the exponential map $\exp:\mathfrak{g}\to G;$ effectively to reach $g=\exp(v)\text{id}$ from identity $\text{id}\in G$, the exponential map dictates that the group element has to move along $v$ for unit time along the manifold. Exponential maps for many groups can be computed analytically, e.g., Rodrigues transformation for rotations, Jordan-Chevalley decomposition 57, or the Cayley Hamilton theorem 58 for matrices. For diffeomorphisms, the Lie algebra is the space of all smooth velocity fields $v:\Omega\to\mathbb{R}^{d}$. There exist iterative methods to approximate the exponential map called the scaling-and-squaring approach 17, 30 which uses the identity
@@
This can be thought of as a special case of  Eq. 6 with $n=2^{M},\epsilon=\frac{1}{n}$ and $u_{1}=\ldots=u_{n}=v$.
By virtue of the exponential map, we can solve the registration problem of finding $\varphi\in G$ by directly optimizing over the Lie algebra $v$. This is because the Lie algebra is a vector space and we can perform, for example, standard Euclidean gradient descent for registration 59, 60, 61. Such methods are called stationary velocity field or shooting methods. At each iteration, one uses the exponential map to get the transformation $\varphi$ from the velocity field $v$, computes the gradient of the registration objective with respect to $\varphi$, pulls back this gradient into the tangent space where $v$ lies
$\nabla_{v}L=\frac{\partial\varphi}{\partial v}\nabla_{\varphi}L$
and finally makes an update to $v$. Traditional methods like DARTEL 17 implement this approach. This is also very commonly used by deep learning methods for registration 30, 62, 63 due to its simplicity. Geodesic shooting methods are more sophisticated implementations of this approach where $\varphi$ is the solution of a time-dependent velocity that follows the geodesic equation; the geodesic is completely determined by the initial velocity $v_{0}\in\mathfrak{g}$.
Adaptive optimization algorithms can be applied to the Lie algebra since it a Euclidean vector space $\mathfrak{g}$. However, there a number of challenges with this method. First, this method requires computing the exponential map and its derivative, both of which need to be iteratively evaluated at each step of gradient descent. This is evident in  Fig. 2 where direct optimization with ANTs runs much faster than the Lie-algebra counterpart. Moreover, the exponential map is only locally diffeomorphic, meaning it is suitable for modelling deviations close to the identity but not for large deformations – this leads to less expressivity and poor performance. In Fig. 2, the greedy SyN method which employs direct optimization significantly outperforms the Lie algebra-based DARTEL. In Fig. 15 we observed that across a large variety of hyper-parameters evaluated via grid search, direct optimization consistently led to better target overlap compared to its Lie algebra counterpart on the LPBA40 dataset. Therefore, we do not consider this method in our work.
4.3.2 Riemannian gradient descent
Solving the registration problem directly on the space of diffeomorphisms avoids repeated computations to and fro via the exponential map. The downside however is that one now has to explicitly account for the curvature and tangent spaces of the manifold. The updates for Riemannian gradient descent 64 at the $t^{\text{th}}$ iteration are
@@ 4.3.2 Riemannian gradient descent
where $\text{where }\nabla_{\varphi}L$ $=\mathbf{g}^{-1}_{\varphi_{t}}\frac{\partial L}{\partial\varphi},$
@@
where one pulls back the Euclidean gradient $\frac{\partial L}{\partial\varphi}$ onto the manifold using the inverse metric tensor $\mathbf{g}$ (which makes the gradient invariant to the parameterization of the manifold of diffeomorphisms) before projecting it to the tangent space using $\text{Proj}_{\varphi_{t}}$. Since the tangent space is a local first-order approximation of the manifold’s surface, we can move along this descent direction by a step-size $\eta$ and compute the updated diffeomorphism $\varphi_{t+1}$, represented as the exponential map from $\varphi_{t}$ computed in the direction of $-\text{Proj}_{\varphi_{t}}(\nabla_{\varphi}L)$.
However, there are a few challenges in optimizing diffeomorphisms using Riemannian gradient descent. First, adaptive optimization algorithms such as RMSProp 25, Adagrad 65 and Adam 26 have become popular because they can handle poorly conditioned optimization problems in deep learning. Variants for optimization on low-dimensional Riemannian manifold exist 66, 67, 68, 69. In contrast to these manifolds, diffeomorphisms are a high-dimensional variable-sized group (e.g., the parameterization of the warp field scales with that of the image size). Therefore, operations like computing the Riemannian metric tensor, and parallel transport of the optimization state variables (momentum and curvature) are very computationally expensive. For diffeomorphisms, computing the parallel transport requires solving a system of partial differential equations, which is computationally expensive. For these reasons, we do not consider direct Riemannian optimization for diffeomorphisms in our work.
4.4 Exploiting the group structure of diffeomorphisms
Diffeomorphisms are imbued with additional structure compared to a Riemannian manifold – they are a Lie group as well. Not all Riemannian manifolds are Lie groups - notable examples of non-Lie group Riemannanian manifolds include the sphere $\mathbb{S}^{n}$, fixed-rank matrices, and the Stiefel and Oblique manifolds 64. The additional Lie group structure of $G(\Omega,\mathbb{R}^{d})$ allows us to exploit the group action to define a gradient descent algorithm that eliminates computationally expensive steps. In the following text, we provide a novel method to compute a descent direction in the group of diffeomorphisms that is computationally efficient and can be used with adaptive optimization algorithms.
Minimizing the Eulerian differential.
Consider a function $U:G\rightarrow\mathbb{R}$ that we aim to minimize. Let $V$ be an admissible Hilbert space of vector fields on $\Omega$ embedded in $C^{1}_{0}(\Omega,\mathbb{R}^{d})$. We define an Eulerian differential in $V$ if there exists a linear form $\partial\bar{U}\in V^{*}$ such that for all $v\in V$:
$\left(\partial\bar{U}(\varphi)|v\right)_{E}=\partial_{t}U(\varphi\circ\varphi^{v}_{0t})\Big{|}_{t=0}$ (8)
This definition of Eulerian differential is different from the one in  54 to perform all updates ($v$) in the tangent space at identity and leverage Jacobian-free descent (see later). The goal is to choose a suitable $v$ such that the directional change of the Eulerian differential along $v$ is negative, making $v$ a descent direction. A more familiar rate of change of $U$ along a curve $v$ is given by the Gateaux derivative:
$\left(\frac{\partial U}{\partial\varphi}\Bigg{|}v\right)_{G}=\partial_{t}U(\varphi+tv)\Big{|}_{t=0}$ (9)
The Eulerian differential is closely related to the Gateaux derivative of $U$ at $\varphi$ as:
$\left(\partial\bar{U}(\varphi)|v\right)_{E}=\left(\frac{\partial U}{\partial\varphi}\Bigg{|}J(\varphi)v\right)_{G}$
using chain rule. The right side is further expanded as:
[…]
where $J(\varphi)(x)=J(\varphi(x))$ with slight abuse of notation. We introduce the Gateaux derivative and relate it to the Eulerian derivative because we typically have access to the Gateaux derivative using automatic differentiation tools like PyTorch, but to perform optimization on the group of diffeomorphisms, we need to compute the Eulerian differential.
Choosing
$v_{d}(x)=-J(\varphi(x))^{T}\frac{\partial U}{\partial\varphi}(\varphi)(x)$
gives us:
[…]
This choice of $v_{d}(x)$ is therefore a descent direction for the Eulerian differential of $U$ at $\varphi$. To perform gradient descent on the Eulerian differential at $\varphi$, we need to compute the descent direction $v_{d}$, perform the exponential map with a small learning rate $\eta_{t}$, and perform the update:
$\varphi_{t+1}=\varphi_{t}\circ\exp_{id}(\eta_{t}v_{d})$
For small enough $\eta_{t}$, the exponential map can be approximated with a retraction map (i.e. $\exp_{id}(\eta_{t}v_{d})\approx id+\eta_{t}v_{d}$), which is quick to compute.
We quickly contextualize the key differences between Gateaux gradient descent and our proposed Eulerian descent. First, the steepest descent direction in Gateaux gradient descent is $-\frac{\partial U}{\partial\varphi}(\varphi)$, whereas it is $-J(\varphi)^{T}\frac{\partial U}{\partial\varphi}(\varphi)$ in Eulerian descent. Second, the update rule in Gateaux gradient descent is $\varphi_{t+1}=\varphi_{t}-\eta_{t}\frac{\partial U}{\partial\varphi}(\varphi)$, whereas it is $\varphi_{t+1}=\varphi_{t}\circ\exp_{id}(\eta_{t}v_{d})$ in Eulerian descent. These two differences capture the essence of performing optimization on the group of diffeomorphisms in contrast to optimizing on the (Euclidean) ambient space directly.
Adaptive optimization on diffeomorphisms.
Note that for small enough $t$, the descent direction $v_{d}(x)$ can also be interpreted as a vector in the tangent space at identity, with $\varphi_{0t}^{v}=\exp_{id}(tv)$ since $\varphi_{00}^{v}=id$, and $\partial_{t}\varphi_{0t}^{v}|_{t=0}=v$. Descent directions over gradient descent iterations $i$ denoted as $v_{d}^{(i)}$ all lie on the same vector space, i.e. the tangent space at identity. Therefore, first order algorithms like Adam can be applied on the sequence of descent directions $v_{d}^{(i)}$ which now lie in the same vector space, without requiring computing the metric tensor, parallel transport or change of coordinates (charts) throughout the optimization process. This framework leveraging the group structure forms the core of our adaptive optimization algorithm for diffeomorphisms. Our framework is therefore a significant advantage over Riemannian optimization methods which require parallel transport of the momentum and curvature vectors at each iteration.
4.4.1 Jacobian-Free Eulerian Descent
We provided an obvious choice of descent direction $v_{d}(x)$ for the Eulerian differential of $U$ at $\varphi$. The Gateaux derivative $\frac{\partial U}{\partial\varphi}$ is readily obtained using automatic differentiation tools like PyTorch. However, the descent direction requires us to multiply this derivative with the Jacobian of the diffeomorphism $J(\varphi)$, which may be computationally expensive. However, in most diffeomorphic image registration applications, the role of the diffeomorphism is warp the image by performing local translations, scaling and shearing without introducing large local rotations. Mathematically, we consider the polar form of the Jacobian $J(\varphi)(x)=U(x)P(x)$ where $U(x)$ is a unitary matrix, and $P(x)$ is a positive definite matrix. We assume that for most applications, $U(x)\approx I_{d\times d}$, making $J(\varphi)(x)$ positive definite. With this assumption, we can choose the modified descent direction
$v_{d}^{\prime}(x)=-\frac{\partial U}{\partial\varphi}(\varphi)(x)$
and the Eulerian differential at $\varphi$ is
[…]
since $v_{d}^{\prime}(x)^{T}J(\varphi(x))v_{d}^{\prime}(x)\geq 0$ for all $x\in\Omega$, owing to the (assumed) positive definiteness of $J(\varphi)(x)$. For all experiments, Jacobian-free descent directions $v_{d}^{\prime}(x)$ are used, and they provide faster runtime and with same accuracy. Adaptive first-order optimization can now be performed on this modified sequence on descent directions $v_{d}^{\prime(i)}(x)$, saving significant computational and memory overhead by avoiding computation of $J(\varphi)$.
Note that this algorithm using the Eulerian differential is only possible due to the group structure of diffeomorphisms. For an arbitrary Riemannanian manifold $\mathcal{M}$ and points $\varphi,\varphi_{0t}^{v}\in\mathcal{M}$, the operation $\varphi\circ\varphi_{0t}^{v}$ does not make sense. The additional group structure of $G(\Omega,\mathbb{R}^{d})$ allows us to propose a novel Eulerian descent algorithm without performing Lie algebra optimization, or Riemannian gradient descent, both of which are computationally expensive for diffeomorphisms.
4.5 Interpolation strategies for multi-scale registration